OpenStudy (anonymous):

A customer is quoted a price of $390 for carpet including installation for one bedroom in her house. She wants to use the same carpet in the living room, which has double the length and width of her bedroom.

OpenStudy (anonymous):

Assuming that these rooms are rectangular, Area = length X Width. Therefore \[x \times y = 360\] \[2x \times 2y = 4(x \times y) = 4(390)= 1560\]

OpenStudy (anonymous):

oops, I meant to write 390* for the first equation, sorry